OpenStudy (anonymous):

Simplify: 5 • [15 – (3 + 8)] – (2 • 8) + 9

OpenStudy (anonymous):

First add 3 and 8 which equals?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Now we have 5* (15 - 11) - 2*8 + 9 Yes, then subtract 15 and 11 which equals?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Now we have 5* 4 - 2*8 + 9 First multiply 5 and 4.

OpenStudy (anonymous):

20

OpenStudy (anonymous):

20 - 2 * 8 + 9 Then, multiply 2 and 8 and we get...?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

20 - 16 + 9 Now, subtract 20 and 16.

OpenStudy (anonymous):

16

OpenStudy (anonymous):

i got 5 when i did the problem

OpenStudy (anonymous):

No not quite. How did did you get 5?

OpenStudy (anonymous):

i did 3+8 got 11

OpenStudy (anonymous):

11-15=4

OpenStudy (anonymous):

2*8=16

OpenStudy (anonymous):

5*4=20

OpenStudy (anonymous):

16+9=25

OpenStudy (anonymous):

20-25=5

OpenStudy (anonymous):

You must solve the addition and subtraction from left to right, so then you have to solve the subtraction first for 20 - 16 + 9.

OpenStudy (anonymous):

so its 25

OpenStudy (anonymous):

25? when you subtract 20 and 16 you get 4... when you add 4 and 9 you get 13. Where did you get 25?

OpenStudy (anonymous):

im trippin my bad i was look at the paper next to me and wrote 25 lol

OpenStudy (anonymous):

haha its okay. Just read the question left to right when dealing with multiplication and division or addition and subtraction. (Even parenthesis sometimes).

OpenStudy (anonymous):

ok thanks!! :) :) :)

OpenStudy (anonymous):

No problem!